# Polynomial Time Convergence of the Iterative Evaluation of Datalogo Programs

### 21 Dec 2023 v1

@@ 1.1.2 $\text{\sf Datalog}^{\circ}$
As a classic example of a nonlinear $\text{\sf Datalog}^{\circ}$ program, consider the following alternate program over the tropical semiring for computing the shortest path length $T(X,Y)$ between all pairs $X,Y$ of vertices in a directed graph specified by an edge relation $E(X,Y)$, where the semiring element associated with $E(X,Y)$ is the length of the directed edge $(X,Y)$:
@@
In the $\text{\sf Datalog}^{\circ}$ program in Example 4, if $(a,b)$ was a ground atom in $T$ with associated semiring element $x$, meaning that the shortest known directed path from vertex $a$ to vertex $b$ has length $x$, and $(b,c)$ was a ground atom in $T$ with associated semiring element $y$, meaning that the shortest known directed path from vertex $b$ to vertex $c$ has length $y$, then the ICO would make the semiring element associated with $(a,c)$ the minimum (as minimum is the addition operation in the tropical semiring) of its current value and $x+y$ (as normal additional is the multiplication operation in the tropical semiring).
1.1.3 Stability
Given a semiring $\bm{S}=(S,\oplus,\otimes,0,1)$ and $u\in S$, let $u^{(p)}:=1\oplus u\oplus u^{2}\oplus\cdots\oplus u^{p}$, where $u^{i}:=u\otimes u\otimes\cdots\otimes u$ ($i$ times). Then $u\in S$ is $p$-stable if $u^{(p)}=u^{(p+1)}$, and semiring $\bm{S}$ is $p$-stable if every element $u\in S$ is $p$-stable. A function $f:S^{n}\to S^{n}$ is $p$-stable if $f^{(p+1)}(\bm{0})=f^{(p)}(\bm{0})$, where $\bm{0}$ is the all zero vector, and $f^{(k)}$ is the $k$-fold composition of $f$ with itself. The stability index of $f$ is the smallest $p$ such that $f$ is $p$-stable. The convergence time of a $\text{\sf Datalog}^{\circ}$ program is the stability index of its ICO. It is important to note that a semiring is $p$-stable if and only if every linear function $f:S\to S$ is $p$-stable.
1.1.4 Grounding the ICO
Since the final associated semiring values of the ground atoms in an IDB are not initially known, it is natural to think of them as (IDB) variables. Then the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$program is a map $\bm{f}:S^{n}\rightarrow S^{n}$, where $n$ is the number of ground atoms for the IDB’s that ever have a nonzero value at some point in the iterative evaluation of the $\text{\sf Datalog}^{\circ}$ program. For instance, in (6), there would be one variable for each pair $(x,y)$ of vertices where there is a directed path from $x$ to $y$ in the graph. So the grounded version of the ICO of a $\text{\sf Datalog}^{\circ}$ program has the following form:
$-\displaystyle X_{1}\text{ :- }$- $f_{1}(X_{1},\ldots,X_{n})$
$\ldots$ (8)
$-\displaystyle X_{n}\text{ :- }$- $f_{n}(X_{1},\ldots,X_{n})$
where the $X_{i}$’s are the IDB variables, and $f_{i}$ is the component of $\bm{f}$ corresponding to the IDB variable $X_{i}$. Note that each component function $f_{i}$ is a multivariate polynomial in the IBD variables of degree at most the maximum number of terms in any product in the body of some rule in the $\text{\sf Datalog}^{\circ}$ program. After $q$ iterations of the iterative evaluation of a $\text{\sf Datalog}^{\circ}$ program, the semiring value associated with the ground atom corresponding to $X_{i}$ will be:
${f}_{i}^{(q)}(\bm{0})$ (9)
1.1.5 Context Free Languages
It is convenient to reason about the formal expansion of $f_{i}^{(q)}(\bm{0})$ using context-free languages (CFL). To explain this, it is probably best to start with a concrete example.
@@ 1.1.5 Context Free Languages
$\begin{bmatrix}A\\ B\end{bmatrix}\to\begin{bmatrix}aAB+bB+c\\ cAB+bA+a\end{bmatrix}$ (14)
@@
$f_{i}^{(q)}(0)$ $=\sum_{T\in\mathcal{T}^{i}_{q}}Z(T).$ (15)
That is, the value of the semiring value associated with a ground atom after in a IDB after $q$ iterations is the sum over the product of the leaves of parse trees of depth at most $q$ and rooted at $X_{i}$.
1.1.6 Parikh’s Theorem
The upper bound on the time to convergence for iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs in  [10] essentially relied on black-box application of Parikh’s theorem [19]. The Parikh image of a word $w\in\Sigma^{*}$, denoted by $\Psi(w)$, is the vector $\Psi(w)=(k_{1},\dots,k_{\sigma})\in\mathbb{N}^{\sigma}$ where $k_{i}$ is the number of occurrences of the letter $a_{i}\in\Sigma$ that occur in the word $w$ (So $|\Sigma|=\sigma$). Then using our assumption that the underlying semiring $\bm{S}$ is commutative, we observe that
$f_{i}^{(q)}(\bm{0})$ $=\sum_{T\in\mathcal{T}^{i}_{q}}Z(T)=\sum_{T\in\mathcal{T}^{i}_{q}}\prod_{j=1}^{\sigma}a_{j}^{\Psi_{j}(Y(T))}$ (16)
where $\Psi_{j}(w)$ is component $j$ of the Parikh image. One version of Parikh’s theorem [19] states that the Parikh images of the words in in a context free language forms a semi-linear set. A set is then semi-linear if it is a finite union of linear sets. A set $\mathcal{L}\subseteq\mathbb{N}^{\sigma}$ is said to be linear if there exist offset vector $\bm{v}_{0}$ and basis vectors $\bm{v}_{1},\ldots,\bm{v}_{\ell}\in\mathbb{N}^{\sigma}$ such that $\mathcal{L}$ is the span of of these vectors, that is if:
@@ 1.1.6 Parikh’s Theorem
If a vector
@@
Figure 1: Illustration of a wedge $W$
1.2 Technical Overview
@@
  1. 1.
    A word $w$ is in $L$ if and only if $\Psi(w)$ is in some linear set $\mathcal{L}\in\mathcal{M}$.
  2. 2.
    Every linear set $\mathcal{L}\in\mathcal{M}$ has an associated offset vector $v_{0}$ and basis vectors $v_{1},\ldots v_{\ell}$, with the property that :
@@
      For each vector $\bm{v}=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}$ in the span of $\mathcal{L}$ it is the case that there is a word $w\in L$, with $\Psi(w)=\bm{v}$, where $w$ can be generated by a parse tree with depth at most $(k+1)n(n+3)/2$, where $k=k_{1}+k_{2}+\ldots+k_{\ell}$.
The proof of Theorem 1.2 is given in Section 2. The most important way that Theorem 1.2 extends the standard version of Parikh’s theorem is property 2(d), which upper bounds the depth of some parse tree of a word by the 1-norm of the representation of that word. The bound given in the textbook proof [12] of Parikh’s theorem gives a depth bound that is exponentially large (which essentially yields the exponential upper bound on the convergence time [10]). To achieve property 2(d), our proof contains a constructive forward process $\mathcal{P}$ that creates a linear set $\mathcal{L}\in\mathcal{M}$ from the parse tree $T$ of some word $w\in L$ by removing wedges from $T$. We are careful to design $\mathcal{P}$ so that it is reversible, that is to recover a parse tree from a linear representation we can just reverse the process $\mathcal{P}$. To accomplish this we need that in the forward process every wedge that is removed does not remove any nonterminal from the parse tree. Our process $\mathcal{P}$ insures that the depth of the parse tree for the offset and the wedges for the basis vectors have height $O(n^{2})$.
The second part of the proof of Theorem 1.1 is stated in Lemma 1.3, which states that for every word $w\in L$ there exists a linear set $\mathcal{L}$ in $\mathcal{M}$ such that $w\in\mathcal{L}$ and has a linear representation with respect to the basis vectors of $\mathcal{L}$ that has small support.
@@
Let $L$ be an arbitrary CFL, and $\mathcal{M}$ the semilinear set that is guaranteed to exist in Theorem 1.2. Then for every word $w\in L$ there exists a linear set $\mathcal{L}$ in $\mathcal{M}$ such that $w\in\mathcal{L}$ and $w$ has a linear representation with respect to the basis vectors in $\mathcal{L}$ that has support size at most:
$2(\sigma(n(n+3)/2)(\lg(\lambda+1)+4\sigma\lg\sigma)$
@@
Finally in Section 4 we use Lemma 1.3 to prove Theorem 1.1. The proof considers the context free language $L$ corresponding to a component $i$ of $f$, and then an arbitrary word $w\in L$. By Lemma 1.3 it must be the case that the Parikh image of $w$ has a representation
$\Psi(w)=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}$
in some linear set $\mathcal{L}$ of the semilinear set $\mathcal{M}$ that has small support. If each coefficient $k_{i}$ is small, then by condition 2(d) in Theorem 1.2 it must be the case that there is a low depth parse tree $T$ for $w$, and thus the summand corresponding to $w$ will be added to semiring value associated with $X_{i}$ in a relatively small number of steps in the naive evaluation of the $\text{\sf Datalog}^{\circ}$ program.  Otherwise assume some coordinate, say without loss of generality that $k_{\ell}$ is large. Then note that each of the Parikh images
$\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+(k_{\ell}-j)\bm{v}_{\ell}$
for $j\in[1,p]$ must have been added to the semiring value associated with $X_{i}$ before the summand associated with $w$ was added, and thus by the stability of the product corresponding to $v_{\ell}$, when the Parikh image of $w$ is added to the semiring value associated with $X_{i}$, and thus it must be the case that this semiring value will not change.
2 The Strengthened Parikh’s Theorem
Our goal in this section is to prove Theorem 1.2. A word’s depth is defined as the minimum depth of a parse tree that yields the word. Similarly, we say that a vector $v$’s depth is defined as the minimum depth of a word $w$ with $\Psi(w)=v$.
We begin our proof by defining a semi-linear set $\mathcal{M}$ with the desired properties. Let $c:=n(n+3)/2$ throughout this section. Let $\mathcal{T}^{S}_{c}$ denote the set of all parse trees (starting with the non-terminal $S$) of depth at most $c$. Recall that $T$’s yield, denoted as $Y(T)$, is the word obtained by a parse tree $T$. For a wedge $W$, $Y(W)$ is analogously defined by ignoring the unique non-terminal leaf node in the wedge $W$. Let $N(T)$ denote the set of non-terminals appearing somewhere in $T$. Let $\mathcal{W}^{A}_{c}(T)$ be the collection of wedges that have a non-terminal $A$ as the root, have non-terminals only from $N(T)$ and have depth (or equivalently height) at most $c$. Let $\mathcal{B}_{c}(A,T):=\{Y(W)\;|\;W\in\mathcal{W}^{A}_{c}(T)\}$. For notational brevity, we may use $\mathcal{B}(A,T)$ instead of $\mathcal{B}_{c}(A,T)$.
@@ 2 The Strengthened Parikh’s Theorem
Figure 2: The right tree $T$ is recovered from the left tree $T^{\prime}$ by augmenting the wedge $W$.
@@
Given a parse tree $T$ of depth greater than $c$ starting with $S$, we can obtain a parse tree $T^{\prime}$ starting with $S$ that satisfies the following:
  1. 1.
@@
  2. 2.
    (Reversibility) $T$ can be obtained by replacing one non-terminal leaf $A$ in $T^{\prime}$ with a wedge $W\in\mathcal{W}_{c}^{A}(T^{\prime})$ for some $A\in N(T^{\prime})$.
@@
The first property in the lemma is worth special attention. Suppose we obtained $T^{\prime}$ from $T$ by repeatedly applying the lemma, but without guaranteeing the first property. Suppose $\{W_{i}\}$ are the wedges we removed in the process; so $W_{i}$ is the wedge removed in iteration $i$. Then, we may not be able to augment $T^{\prime}$ with an arbitrary subset of the wedges, which is critical to establish the “if” direction of the first property of Theorem 1.2.
@@
  • The height of $A^{\prime}$ is at most $c$. In other words, the subtree rooted as $A^{\prime}$ has depth at most $c$.
@@
We prove the lemma by an induction on the number of non-terminals. Consider an arbitrary node $v$ of the largest depth, which must be at least $c=(n+3)n/2=(n+1)+n+\ldots+2$. Since the base case $n=1$ is trivial, suppose $n\geq 2$. Consider the unique path from $v$ to the root $S$ in $T$. The height of a node $u$ on the path is defined as the number of nodes below $u$ on the path, including $u$.
By walking up the path, we find the first two occurrences of the same non-terminal $V_{1}$. Say they appear at nodes $v(V_{1})$ and $u(V_{1})$, where $u(V_{1})$ is an ascendant of $v(V_{1})$. Note that $u(V_{1})$ has height at most $n+1$ due to the pigeon hole’s principle. Thus, the two nodes have heights differing by at most $n$.
If $(u(V_{1}),v(V_{1}))$ is a good pair, we are done. If not, it means that the wedge $W(u(V_{1}),v(V_{1}))$ must include a non-terminal that doesn’t appear anywhere else in the tree. Consequently, we now consider the tree $T_{2}$ with the subtree rooted at $u(V_{1})$ removed. Note that the subtree $T_{2}$ has at most $n-1$ non-terminals and has depth at least $n+(n-1)+\ldots+2$. By induction, this implies that $T_{2}$ must have a good pair. It is easy to see that the pair remains to be good with respect to $T$ as well. ∎
@@
Property (1) “only if” direction. Given a parse tree $T$ for $w\in L$, suppose we obtained a sequence of trees $T_{0}=T,T_{1},\ldots,T_{\eta}$ by repeatedly applying Lemma 2.2, where $T_{\eta}\in\mathcal{T}^{S}_{c}$ and $T_{i}$ is obtained from $T_{i-1}$ by deleting a wedge $W_{i}$ in $\mathcal{W}^{A}_{c}(T)$ for some non-terminal $A$ in $N(T)=N(T_{\eta})$; note $\mathcal{W}^{A}_{c}(T)=\mathcal{W}^{A}_{c}(T_{\eta})$ since $N(T)=N(T_{1})=\ldots=N(T_{\eta})$. Let $b_{i}=\Psi(Y(W_{i}))$. Clearly, $\Psi(w)$ can be expressed as $\Psi(Y(T_{\eta}))+\sum_{i=1}^{\eta}b_{i}$.
Since we created a linear set for each parse tree in $\mathcal{T}^{S}_{c}$, thus for $T_{\eta}$, this is a linear representation within $\mathcal{L}$ which consists of offset vector $\Psi(Y(T_{\eta}))$ and basis vectors $\cup_{V\in N(T_{\eta})}\Psi(\mathcal{B}(V,T_{\eta}))$. This proves the “only if” part of the first property.
Property (1) “if” direction and Property (2)(d). Conversely, suppose $w$ has a linear representation within some $\mathcal{L}\in\mathcal{M}$. Say the linear representation is $\Psi(Y(T^{\prime}))+\sum_{i=1}^{k}b_{i}$ for some $T^{\prime}\in\mathcal{T}_{short}$. Note that for each basis vector $b_{i}$, there exists $V_{i}\in N(T^{\prime})$ such that $b_{i}\in\Psi(\mathcal{B}(V_{i},T^{\prime}))$. Because of the way we defined linear sets, $b_{i}=\Psi(Y(W_{i}))$ for some $W_{i}\in\mathcal{W}_{c}^{V_{i}}(T^{\prime})$. We can augment $T^{\prime}$ with the wedges $W_{i}$ in an arbitrary order. Adding each wedge $W_{i}$ increases the tree depth by at most $c$. By repeating this for each $b_{i}$, we obtain a parse tree corresponding to the given linear representation and its depth is at most $c(k+1)$.
@@
3 Small Support Representations
In this section we prove Lemma 1.3. Recall that this lemma states that the Parikh image of every word in a context-free language $L$ has a linear representation with small support.

### 21 Feb 2024 v2

@@ 2.4. Grounding the ICO
Consider the binary recursive formulation in (6), written over a generic semiring.
@@
Suppose $\textsf{ADom}=\{1,2,3,4\}$, and the input EDB contains ground EDB atoms $E(1,2),E(2,3),E(3,4)$, with corresponding (constant) semiring values $e_{12},e_{23},e_{34}$. Then there will be 16 equations and 16 variables in the grounded ICO; For each $a,b\in\{1,\ldots,4\}$ there will a variable $X_{ab}$, and an equation of the form:
[…]
But as many of these variables will always be 00; they are “inactive” and thus they can effectively be ignored from the grounded ICO formulation. Thus effectively one can think of the grounded ICO as having the following $6$ variables and 6 equations:
[…]
2.5. Context Free Languages
It is convenient to reason about the formal expansion of $f_{i}^{(q)}(\bm{0})$ using context-free languages (CFL). See [9] for an introduction to CFLs. To explain this, it is probably best to start with a concrete example.
@@ 2.5. Context Free Languages
(14) $\begin{bmatrix}A\\ B\end{bmatrix}\to\begin{bmatrix}aAB+bB+c\\ cAB+bA+a\end{bmatrix}$
@@
(15) $f_{i}^{(q)}(0)$ $={\bigoplus}_{T\in\mathcal{T}^{i}_{q}}Z(T).$
That is, the value of the semiring value associated with a IDB variable $X_{i}$ after $q$ iterations is the sum of the product of the leaves of parse trees of depth at most $q$ and rooted at $X_{i}$.
2.6. Parikh’s Theorem
The upper bound on the time to convergence for iterative evaluation of $\text{\sf Datalog}^{\circ}$ programs in  [7] essentially relied on black-box application of Parikh’s theorem [12]. See [8] for an introduction to Parikh’s theorem. The Parikh image of a word $w\in\Sigma^{*}$, denoted by $\Psi(w)$, is the vector $\Psi(w)=(k_{1},\dots,k_{\sigma})\in\mathbb{N}^{\sigma}$ where $k_{i}$ is the number of occurrences of the letter $a_{i}\in\Sigma$ that occur in the word $w$ (So $|\Sigma|=\sigma$). Similarly, for a language $L$, define $\Psi(L):=\{\Psi(w)\;|\;w\in L\}$. Then using our assumption that the underlying semiring $\bm{S}$ is commutative, we observe that
(16) $f_{i}^{(q)}(\bm{0})$ $={\bigoplus}_{T\in\mathcal{T}^{i}_{q}}~{}Z(T)={\bigoplus}_{T\in\mathcal{T}^{i}_{q}}~{}{\bigotimes}_{j=1}^{\sigma}~{}a_{j}^{\Psi_{j}(Y(T))}$
where $\Psi_{j}(w)$ is component $j$ of the Parikh image. One version of Parikh’s theorem [12] states that the Parikh images of the words in a context free language forms a semi-linear set. A set is then semi-linear if it is a finite union of linear sets. A set $\mathcal{L}\subseteq\mathbb{N}^{\sigma}$ is said to be linear if there exist offset vector $\bm{v}_{0}$ and basis vectors $\bm{v}_{1},\ldots,\bm{v}_{\ell}\in\mathbb{N}^{\sigma}$ such that $\mathcal{L}$ is the span of these vectors, that is if:
@@ 2.6. Parikh’s Theorem
Here we assume that $0\in\mathbb{N}$. If a vector
@@
Figure 1. Illustration of a wedge $W$
3. Technical Overview
@@
  1. (1)
    $\mathcal{M}=\Psi(L)$.
  2. (2)
    Every linear set $\mathcal{L}\in\mathcal{M}$ has an associated offset vector $v_{0}$ and basis vectors $v_{1},\ldots v_{\ell}$, with the properties that :
@@
      For each vector $\bm{v}=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}$ in the span of $\mathcal{L}\in\mathcal{M}$, there is a word $w\in L$, with $\Psi(w)=\bm{v}$, where $w$ can be generated by a parse tree with depth at most $(k+1)n(n+3)/2$, where $k=k_{1}+k_{2}+\ldots+k_{\ell}$.
  3. (3)
    For any parse tree $T$ of depth $d$ such that $Y(T)\in L$, there exists a linear representation of $\Psi(Y(T))=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{\ell}\bm{v}_{\ell}$ such that $1+k_{1}+k_{2}+\ldots+k_{\ell}\geq d/(n(n+3)/2)$.
The proof of Theorem 3.1 is given in Section 4. The most important way that Theorem 3.1 extends the standard version of Parikh’s theorem is property 2(d), which upper bounds the depth of some parse tree of a word by the 1-norm of the representation of that word. The bound given in the textbook proof [9] of Parikh’s theorem gives a depth bound that is exponentially large. To achieve property 2(d), our proof contains a constructive forward process $\mathcal{P}$ that creates a linear set $\mathcal{L}\in\mathcal{M}$ from the parse tree $T$ of some word $w\in L$ by removing wedges from $T$. We are careful to design $\mathcal{P}$ so that it is reversible; that is, to recover a parse tree from a linear representation we can just reverse the process $\mathcal{P}$. To accomplish this we need that in the forward process every wedge that is removed does not remove any nonterminal from the parse tree. Our process $\mathcal{P}$ ensures that the parse tree for the offset and the wedges for the basis vectors have depth $O(n^{2})$.
The second part of the proof of Theorem 1.1 is stated in Lemma 3.2, which states that for every word $w\in L$ there exists a linear set $\mathcal{L}$ in $\mathcal{M}$ such that $\Psi(w)\in\mathcal{L}$ and $\Psi(w)$ has a linear representation with respect to the basis vectors of $\mathcal{L}$ that has small support.
@@
Let $L$ be an arbitrary context free language. Let $h=2(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma)$. Let $\mathcal{M}$ be the semilinear set that is guaranteed to exist in Theorem 3.1. Let $w$ be a word in $L$. Let $\mathcal{L}$ in $\mathcal{M}$ be a linear set such that $\Psi(w)\in\mathcal{L}$. Let $\Psi(w)=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{m}\bm{v}_{m}$ be a linear representation of $\Psi(w)$ with respect to the offset and basis vectors of $\mathcal{L}$. Then there exists another linear representation $\Psi(w)=\bm{v}_{0}+k^{\prime}_{1}\bm{v}^{\prime}_{1}+\ldots+k^{\prime}_{h}\bm{v}^{\prime}_{h}$ of $\Psi(w)$ with respect to $h$ basis vectors of $\mathcal{L}$ such that $\sum_{i=1}^{m}k_{i}=\sum_{i=1}^{h}k^{\prime}_{i}$.
@@
Finally in Section 6 we use Lemma 3.2 to prove Theorem 1.1. To show that for sufficiently large $q$ it is the case that $f_{i}^{(q)}(\bm{0})=f_{i}^{(q+1)}(\bm{0})$, we show that for every tree $T^{\prime}\in\mathcal{T}^{i}_{q+1}\setminus\mathcal{T}^{i}_{q}$, the corresponding summand $Z(T^{\prime})$ is “absorbed” by the earlier terms, that is:
(17) $\bigoplus_{T\in\mathcal{T}^{i}_{q}}Z(T)\oplus Z(T^{\prime})=\bigoplus_{T\in\mathcal{T}^{i}_{q}}Z(T)$
From property (3) of Theorem 3.1 we know that that there exists a linear representation of $\Psi(Y(T^{\prime}))$ that has a large 1-norm; and from Lemma 3.2 we know that there is a linear representation with the same large 1-norm of $\Psi(Y(T^{\prime}))$ with small support. Thus we can conclude by the pigeonhole principle that one of the basis vectors, in this small support linear representation, must have a coefficient greater than the stability $p$ of the ground set. Finally, Eqn. (17) follows from the stability of the semiring element that corresponds to the semiring element that is the “product” of that basis vector.
4. The Strengthened Parikh’s Theorem
Our goal in this section is to prove Theorem 3.1.
We begin our proof by defining a semi-linear set $\mathcal{M}$ with the desired properties. Let $c:=n(n+3)/2$ throughout this section. Let $\mathcal{T}^{S}_{c}$ denote the set of all parse trees (starting with the non-terminal $S$) of depth at most $c$. Recall that $T$’s yield, denoted as $Y(T)$, is the word obtained by a parse tree $T$. For a wedge $W$, $Y(W)$ is analogously defined by ignoring the unique non-terminal leaf node in the wedge $W$. Let $N(T)$ denote the set of non-terminals that appear in $T$. Let $\mathcal{W}^{A}_{c}(T)$ be the collection of wedges that appear in $T$, have a non-terminal $A$ as the root, and have depth (or equivalently height) at most $c$. Let $\mathcal{B}_{c}(A,T):=\{Y(W)\;|\;W\in\mathcal{W}^{A}_{c}(T)\}$. [^4] While we use notations $\mathcal{W}^{A}_{c}(T)$ and $\mathcal{B}_{c}(A,T)$ for notational brevity, their dependence is on $N(T)$ rather than $T$. For notational brevity, we may use $\mathcal{B}(A,T)$ instead of $\mathcal{B}_{c}(A,T)$.
@@ 4. The Strengthened Parikh’s Theorem
Figure 2. The right tree $T$ is recovered from the left tree $T^{\prime}$ by augmenting the wedge $W$.
@@
Given a parse tree $T$ of depth greater than $c$ starting with non-terminal $S$, we can obtain a parse tree $T^{\prime}$ starting with $S$ that satisfies the following:
  1. (1)
@@
  2. (2)
    (Reversibility) $T$ can be obtained by replacing one non-terminal $A$ in $T^{\prime}$ with a wedge $W\in\mathcal{W}_{c}^{A}(T^{\prime})$ for some $A\in N(T^{\prime})$.
@@
The first property in the lemma is worth special attention. Suppose we obtained $T^{\prime}$ from $T$ by repeatedly applying the lemma, but without guaranteeing the first property. Suppose $\{W_{1},W_{2},\dots\}$ are the wedges we removed in the process; so $W_{i}$ is the wedge removed in iteration $i$. Then, we may not be able to augment $T^{\prime}$ with an arbitrary subset of the wedges, which is critical to establish the $\Psi(L)\supseteq\mathcal{M}$ direction of the first property of Theorem 3.1.
@@
  • The height of $A^{\prime}$ is at most $c$. In other words, the subtree rooted at $A^{\prime}$ has depth at most $c$.
@@
We prove the lemma by an induction on the number of non-terminals. Consider an arbitrary node $v$ of the largest depth, which must be at least $c=(n+3)n/2\geq(n+1)+n+\ldots+2$. Since the base case $n=1$ is trivial, suppose $n\geq 2$. Consider the unique path from $v$ to the root non-terminal $S$ in $T$. The height of a node $u$ on the path is defined as the number of nodes below $u$ on the path, including $u$.
Consider walking from $u$ towards the root. On this path, consider the first time two occurrences of the same non-terminal $V_{1}$ appear. Say they appear at nodes $v(V_{1})$ and $u(V_{1})$, where $u(V_{1})$ is an ascendant of $v(V_{1})$. Observe that $u(V_{1})$ has height at most $n+1$ due to the pigeon hole principle. This is because some non-terminal must repeat among $n+1$ nodes.
If $(u(V_{1}),v(V_{1}))$ is a good pair, we are done. If not, it means that the wedge $W(u(V_{1}),v(V_{1}))$ must include a non-terminal that doesn’t appear anywhere else in the tree. Consider the tree $T_{2}$ with the subtree rooted at $u(V_{1})$ removed. This subtree $T_{2}$ has at most $n-1$ non-terminals and has depth at least $n+(n-1)+\ldots+2$. By induction, this implies that $T_{2}$ must have a good pair $(u_{2},v_{2})$.
Finally, it is easy to see that the pair remains to be good with respect to $T$ as well: First, $u_{2}$ has height at most $(n+1)+(n+(n-1)+\ldots+2)=c$ in the tree $T$. Second, the wedge indexed by $(u_{2},v_{2})$ does not intersect the subtree rooted at $u(V_{1})$, and therefore, the set of non-terminals remains unchanged after removing the wedge from $T$, just as it does when we remove the wedge from $T_{2}$. ∎
@@
Property (1) $\mathcal{M}\supseteq\Psi(L)$ and Property (3). Given a parse tree $T$ for $w\in L$, suppose we obtained a sequence of trees $T_{0}=T,T_{1},\ldots,T_{\eta}$ by repeatedly applying Lemma 4.2, where $T_{\eta}\in\mathcal{T}^{S}_{c}$ and $T_{i}$ is obtained from $T_{i-1}$ by deleting a wedge $W_{i}$ in $\mathcal{W}^{A}_{c}(T)$ for some non-terminal $A$ in $N(T)=N(T_{\eta})$; note $\mathcal{W}^{A}_{c}(T)=\mathcal{W}^{A}_{c}(T_{\eta})$ since $N(T)=N(T_{1})=\ldots=N(T_{\eta})$. Let $\bm{b}_{i}=\Psi(Y(W_{i}))$. Clearly, $\Psi(w)$ can be expressed as $\Psi(Y(T_{\eta}))+\sum_{i=1}^{\eta}\bm{b}_{i}$.
Since we created a linear set for each parse tree in $\mathcal{T}^{S}_{c}$, thus for $T_{\eta}$, this is a linear representation within $\mathcal{L}$ which consists of offset vector $\Psi(Y(T_{\eta}))$ and basis vectors $\cup_{V\in N(T_{\eta})}\Psi(\mathcal{B}(V,T_{\eta}))$. This proves $\mathcal{M}\supseteq\Psi(L)$ part of the first property.
Property (3) is immediate from the above: Suppose that the parse tree $T$ has depth $d$. The two trees $T_{i-1}$ and $T_{i}$ have depths differing by at most $c$ since we obtained $T_{i}$ from $T_{i-1}$ by deleting a wedge of depth at most $c$. Further, the last tree $T_{\eta}$ has depth at most $c$ as well. Thus, $c(\eta+1)\geq d$. Since $\eta=k$ where $k$ is described as in the theorem, we have proven this property.
Property (1) $\mathcal{M}\subseteq\Psi(L)$ and Property (2)(d). Conversely, suppose $w$ has a linear representation within some $\mathcal{L}\in\mathcal{M}$. Say the linear representation is $\Psi(Y(T^{\prime}))+\sum_{i=1}^{k}\bm{b}_{i}$ for some $T^{\prime}\in\mathcal{T}_{c}^{S}$. Note that for each basis vector $\bm{b}_{i}$, there exists $V_{i}\in N(T^{\prime})$ such that $\bm{b}_{i}\in\Psi(\mathcal{B}(V_{i},T^{\prime}))$. Because of the way we defined linear sets, $\bm{b}_{i}=\Psi(Y(W_{i}))$ for some $W_{i}\in\mathcal{W}_{c}^{V_{i}}(T^{\prime})$. We can augment $T^{\prime}$ with the wedges $W_{i}$ in an arbitrary order. Adding each wedge $W_{i}$ increases the tree depth by at most $c$. By repeating this for each $\bm{b}_{i}$, we obtain a parse tree $T$ such that $\Psi(Y(T))=\Psi(w)$. This proves $\mathcal{M}\subseteq\Psi(L)$ part of Property (1). Furthermore, Property (2)(d) follows since $T$ has depth at most at most $c(k+1)$.
@@
Together, the proof of the above properties prove Theorem 3.1.
5. Small Support Representations
In this section we prove Lemma 3.2. The lemma shows that any linear representation of the Parikh image of a word in a context-free language $L$ can be converted into another linear representation with equal 1-norm, but with small support.
Lemma 5.1.
[Lemma 3.2 Restated] Let $L$ be an arbitrary context free language. Let $h:=2(\sigma(n(n+3)/2)\lg(\lambda+1)+4\sigma\lg\sigma)$. Let $\mathcal{M}$ be the semilinear set that is guaranteed to exist in Theorem 3.1. Let $w$ be a word in $L$. Let $\mathcal{L}$ in $\mathcal{M}$ be a linear set such that $\Psi(w)\in\mathcal{L}$. Let $\Psi(w)=\bm{v}_{0}+k_{1}\bm{v}_{1}+\ldots+k_{m}\bm{v}_{m}$ be a linear representation of $\Psi(w)$ with respect to the offset and basis vectors of $\mathcal{L}$. Then there exists another linear representation $\Psi(w)=\bm{v}_{0}+k^{\prime}_{1}\bm{v}^{\prime}_{1}+\ldots+k^{\prime}_{h}\bm{v}^{\prime}_{h}$ of $\Psi(w)$ with respect to $h$ basis vectors of $\mathcal{L}$ such that $\sum_{i=1}^{m}k_{i}=\sum_{i=1}^{h}k^{\prime}_{i}$.